Calculate the exponential distribution, expected value, and variance of the given data.

μ = 2, x1 = 4, x2 = 1

1
Expert's answer
2021-06-14T17:05:44-0400

"f(x,\mu)=x = \begin{cases}
 \mu e^{-\mu x}&\text{if } x\ge0 \\
 0 &\text{if } x<0
\end{cases}"


"P(1<x<4)=\int^4_12e^{-4x}dx=-\frac{1}{2}(e^{-16}-e^{-4})=0.00916"


"E(X)=1/\mu=1/2"


"V(X)=1/\mu^2=1/4"